# Evolution of Semi-convective Staircases in Rotating Flows: Consequences for Fuzzy Cores in Giant Planets

### v1

@@ 4 Discussion
$\dfrac{D{\mbox{{\boldmath$u$}}}}{Dt}+\dfrac{\rm Pr}{\rm Ek}{\mbox{{\boldmath$\hat{z}$}}}\times{\mbox{{\boldmath$u$}}}=-\nabla P^{\prime}+\mathrm{Ra}\mathrm{Pr}\left(R_{\rho}C^{\prime}-T^{\prime}\right){\mbox{{\boldmath$\hat{z}$}}}+\mathrm{Pr}\nabla^{2}{\mbox{{\boldmath$u$}}}\,,$ (2)
$\dfrac{DC}{Dt}=\mathrm{Le}^{-1}\nabla^{2}C\,,$ (3)
$\dfrac{DT}{Dt}=\nabla^{2}T\,,$ (4)
[…]
In the equations above, there are 5 dimensionless numbers that characterize the evolution of the flow. These are the Rayleigh number Ra, Ekman number Ek, density ratio $R_{\rho}$, Prandtl number Pr, and Lewis number Le, which are defined respectively as
$\mathrm{Ra}=\frac{\alpha g|\partial_{z}T_{0}|L^{4}_{z}}{\kappa_{T}\nu},\hskip 2.84544pt\mathrm{Ek}=\frac{\nu}{2\Omega L^{2}_{z}},$ (5)
$R_{\rho}=\frac{\beta|\partial_{z}C_{0}|}{\alpha|\partial_{z}T_{0}|},\hskip 2.84544pt\mathrm{Pr}=\frac{\nu}{\kappa_{T}},\hskip 2.84544pt\mathrm{Le}=\frac{\kappa_{T}}{\kappa_{C}},$ (6)
where $\Omega$ is the rotation rate, $\nu$ is the kinematic viscosity, and $\kappa_{T}$ and $\kappa_{C}$ are the thermal and compositional diffusivities, respectively. Note that $R_{\rho}$ is the ratio between the stabilizing and destabilizing effect of the compositional and thermal buoyancy.
The initial structure of the fluid is composed of a single layer, where the temperature and composition are respectively
$T_{0}(z)=1-z\,,\qquad C_{0}(z)=1-z~{}.$ (7)
We fix the flux of temperature and composition at the top and bottom of the domain, so the boundary conditions for temperature and composition are $\partial_{z}T|_{z=0,1}=\partial_{z}C|_{z=0,1}=-1$ The boundary conditions for the velocity are impenetrable and stress free at both boundaries ($\hat{z}\cdot{\mbox{{\boldmath$u$}}}=\hat{x}\cdot\partial_{z}{\mbox{{\boldmath$u$}}}=\hat{y}\cdot\partial_{z}{\mbox{{\boldmath$u$}}}=0$ at $z=0,1$).
@@
We time-evolve equations A1–A4 using the Dedalus pseudospectral code (Burns et al., 2020), using timestepper SBDF2 (Wang & Ruuth, 2008) and CFL safety factor 0.3. All variables are represented using a Chebyshev series with 384 terms for $z\in[0,1]$ and Fourier series with 384 terms in the periodic $x$ and $y$ directions, where $x\in[0,0.75]$, $y\in[0,0.75]$. To avoid aliasing errors, we use the 3/2-dealiasing rule in all directions.

### v3

@@ Appendix A Fluid equations and numerical methods
$\dfrac{D{\mbox{{\boldmath$u$}}}}{Dt}+\dfrac{\rm Pr}{\rm Ek}{\mbox{{\boldmath$\hat{z}$}}}\times{\mbox{{\boldmath$u$}}}=-\nabla P^{\prime}+\mathrm{Ra}\mathrm{Pr}\left(R_{\rho}C^{\prime}-T^{\prime}\right){\mbox{{\boldmath$\hat{z}$}}}+\mathrm{Pr}\nabla^{2}{\mbox{{\boldmath$u$}}}\,,$ (A2)
$\dfrac{DC}{Dt}=\mathrm{Le}^{-1}\nabla^{2}C\,,$ (A3)
$\dfrac{DT}{Dt}=\nabla^{2}T\,,$ (A4)
where $D/Dt=\partial_{t}+{\mbox{{\boldmath$u$}}}\cdot{\mbox{{\boldmath$\nabla$}}}$, and $u$ is the velocity field.
There are 5 dimensionless numbers that characterize the evolution of the flow. These are the Rayleigh number Ra, Ekman number Ek, density ratio $R_{\rho}$, Prandtl number Pr, and Lewis number Le, which are defined respectively as
$\mathrm{Ra}=\frac{\alpha g|\partial_{z}T_{0}|L^{4}_{z}}{\kappa_{T}\nu},\hskip 2.84544pt\mathrm{Ek}=\frac{\nu}{2\Omega L^{2}_{z}},$ (A5)
$R_{\rho}=\frac{\beta|\partial_{z}C_{0}|}{\alpha|\partial_{z}T_{0}|},\hskip 2.84544pt\mathrm{Pr}=\frac{\nu}{\kappa_{T}},\hskip 2.84544pt\mathrm{Le}=\frac{\kappa_{T}}{\kappa_{C}},$ (A6)
where $\Omega$ is the rotation rate, $\nu$ is the kinematic viscosity, and $\kappa_{T}$ and $\kappa_{C}$ are the thermal and compositional diffusivities, respectively. $R_{\rho}$ is the ratio of the stabilizing and destabilizing effect of the compositional and thermal buoyancy.
We initialize the fluid with linear distributions of temperature and composition, $T_{0}(z)=1-z,\quad C_{0}(z)=1-z$, and fix the flux of temperature and composition at the top and bottom of the domain, so the boundary conditions for temperature and composition are $\partial_{z}T|_{z=0,1}=\partial_{z}C|_{z=0,1}=-1$ The boundary conditions for the velocity are impenetrable and stress free at both boundaries ($\hat{z}\cdot{\mbox{{\boldmath$u$}}}=\hat{x}\cdot\partial_{z}{\mbox{{\boldmath$u$}}}=\hat{y}\cdot\partial_{z}{\mbox{{\boldmath$u$}}}=0$ at $z=0,1$).
@@
We time-evolve equations A1–A4 using the Dedalus pseudospectral solver (Burns et al., 2020) version 3, using timestepper SBDF2 (Wang & Ruuth, 2008) and CFL safety factor 0.2. All variables are represented using a Chebyshev series with 384 terms for $z\in[0,1]$ and Fourier series with 384 terms in the periodic $x$ and $y$ directions, where $x,~{}y\in[0,0.75]$. We use the 3/2-dealiasing rule in all directions, so that nonlinearities are calculated in physical space on a $576^{3}$ grid. To start our simulations, we add random-noise temperature perturbations sampled from a normal distribution with a magnitude of $10^{-5}$ compared to the initial temperature field. Using 2048 cores, the non-rotating and rotating runs took 20 and 55 days to finish, respectively.